OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

answer my question

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2